OpenStudy (anonymous):

4z^2+16z+15/2z^2-z-15

OpenStudy (anonymous):

???

OpenStudy (anonymous):

1st half is on top last half is on bottom

OpenStudy (anonymous):

simplify???

OpenStudy (anonymous):

i got it never mind

OpenStudy (anonymous):

2-16-1 = -15

OpenStudy (callisto):

\[\frac{4z^2+16z+15}{2z^2-z-15}=\frac{(2x+3)(2z+5)}{(2z+5)(x-3)}\]Cancel the common factor, what do you get?

OpenStudy (anonymous):

thanks man thats even better

OpenStudy (callisto):

welcome!~